OpenStudy (anonymous):

let log A=3; log C=2; log D=5 b b b what is the value of log base b of D^2/C^3A? A. 1 B. 4.44 C. 19 D. 0.549

OpenStudy (mathmale):

First of all, using parentheses would make it easier for your readers to grasp exactly what you mean. Instead of D^2/c^3a, I'd write the following for clarity: D^2/(A*C^3), or, better yet, \[\frac{ D ^{2} }{ AC ^{3} }.\] Applying the appropriate rules of logs, \[\log _{b}\left( \frac{ D ^{2} }{ AC ^{3} } \right)=\log _{b}D ^{2}-\log _{b}(AC ^{3}).\] Can you finish this problem? If not, what do you need to know in order to finish it?

OpenStudy (anonymous):

i think i can finish it thank you

OpenStudy (mathmale):

You're very welcome, Katie. Happy New Year!

OpenStudy (anonymous):

Happy New Year to you too!